OpenStudy (osanseviero):

When an equation has no solution, when is it an identity and when does it has infinite solutions?

OpenStudy (anonymous):

an equation that has no solution is a false statement. for example: x^2 = -5 is a false statement (if we restrict our range to real numbers) an identity is an equation that reduces to True, or when something equals itself. for example: x^2 = 2x^2/2. also all trig identities are of this nature. an equation with infinite solutions is an equation that is not an identity, but contains an infinite number of solutions. for example. y = x^2. there are an infinite number of solutions that make this statement true

OpenStudy (osanseviero):

Thanks :D